OpenStudy (anonymous):

If the point (1, 5) lies on a circle whose center is at the origin, what is the equation of the circle? x2 + y2 = 26 x2 + y2 = 16 x2 + y2 = 13 x2 + y2 = 11

OpenStudy (hoblos):

r^2 = 5^2 +1 =26 so x2 + y2 = 26

OpenStudy (experimentx):

the first one ... as it satisfies the point on its eqn